# Random Forest for Dynamic Risk Prediction of Recurrent Events: A Pseudo-Observation Approach

### v2

@@ 2 Notation and Censored Longitudinal Data Structure
Our prediction target can now be formally expressed as $P(T_{i}(t)\geq\tau|Z_{i}(t))$, i.e., at a pre-specified time point $t$, the probability of remaining recurrent-event-free over a subsequent follow-up period of length $\tau$ given covariates $Z_{i}(t)$. In the following, we will occasionally suppress the subscript $i$ and write $P(T(t)\geq\tau|Z(t))$ for ease of presentation.
Figure 1 uses three hypothetical subjects to illustrate the above three steps of converting traditional recurrent event data $\{X^{*}_{i,j},\delta_{i,j}\}$ into the censored longitudinal data format $\{X_{i}(t),\delta_{i}(t):t\in\mathcal{T}_{i}\}$. In Section 4, we will use the constructed censored longitudinal data to create longitudinal pseudo-observations that will be inputs of a random forest algorithm, which we now review.
[…]
3 Random Forest for Uncensored Longitudinal Outcomes: A Review
@@
The number of possible trees is combinatorically large and precludes efficient global optimization (the number of binary trees with $K$ leaves is $(2K-3)!!$!). Greedy algorithms have been developed to approximate the optimal global tree by myopically optimizing prediction error at the start of each branch. We will focus on binary trees in this paper for their popularity and effectiveness. The loss associated with the prediction error for a branch is often termed “impurity” which measures how similarly the observations behave on either side of the binary split. The branching procedure halts, for example, when the number of observations in a terminal node or the number of terminal nodes reaches respective thresholds. Advantages of a regression tree include invariance to monotonic transformation of predictors, flexible approximation to potentially severe nonlinearities, and the capacity to approximate $L-1$ way interactions for a tree with depth $L$.
@@
Let $(Y_{i}(t)$, $\bm{Z}_{i}(t))$ be a pair of a continuous outcome and a set of $p$-dimensional covariates, measured at time $t\in\mathcal{T}_{i}$ for $i=1,2,...,n$; here we use ${\mathcal{T}}_{i}$’s introduced in Section 2 with shared measurement timings to align more closely with our construction in Section 4, although the algorithm presented herein is applicable to scenarios with subject-specific measurement timings. In Section 4, we will substitute $Y_{i}(t)$ here with a longitudinal outcome ($\widehat{S}_{i}^{\tau}(t)$) that is constructed based on pseudo-observation technique to overcome censoring. Also note that the $\bm{Z}_{i}(t)$ may include time $t$. The following algorithm assumes $Y_{i}(t)=g\{\bm{Z}_{i}(t)\}+\epsilon_{i}(t)$ with $\mathbb{E}[\epsilon_{i}(t)]=0$ and is designed to estimate $g(\cdot)$ which describes a potentially complex relationship between the outcome and the predictors.
@@
      (Choose branching): Choosing a covariate among $\bm{W}$ and the split value $w_{kl}$ will result in two daughter nodes, $\{N_{w_{kl},1},N_{w_{kl},0}\}$. Criteria for selecting the best variable and split vary by outcome type and metric of prediction accuracy. For continuous outcomes, the split with minimal sum of squared errors (SSE), is selected, though for other responses types, loss functions such as the Gini Index may be selected. For any particular node split under consideration using covariate $w$, that partitions individuals into two groups ($\{N_{w,0},N_{w,1}\}$) of sizes $n_{w,0}$ and $n_{w,1}$, we minimize the criterion $SSE(w)=\sum_{c=0,1}\sum_{q=1}^{n_{w,c}}(Y_{q}-\overline{Y_{c}})^{2}$, where $\overline{Y_{c}}=\frac{1}{n_{w,c}}\sum_{q=1}^{n_{w,c}}Y_{q,c}=0,1$.
@@
      (Stopping criteria and prediction from the grown tree) Repeat Step 2a and 2b until one of the following hyperparameters has been met: minimum node size ($\min\{n_{w,c},c=0,1\}$) is set to $40$ in this paper) is reached, or $SSE(w)$ does not decrease any further with more branches. This results in a single binary decision tree. For each terminal node (“leaf”) of the tree $C$, predict $Y_{i}(t)$ by the average of the outcome for observations in leaf $C$. Let the prediction be $\widehat{Y}_{b}(\bm{Z}_{i}(t);\widehat{\bm{\theta}}_{b})$, where $\widehat{\bm{\theta}}_{b}$ is the collection of terminal node predictions in the $b$-th tree.
@@
    (Ensemble prediction by random forest): The final prediction for an individual $i^{\prime}$ with covariates $\bm{Z}_{i^{\prime}}(t)$ is given by the average of all the outputs from $B$ binary decision trees above: $\widehat{Y}_{i^{\prime}}(t)=\frac{1}{B}\sum_{b=1}^{B}\widehat{Y}_{b}(\bm{Z}_{i^{\prime}}(t);\widehat{\bm{\theta}}_{b})$.
@@
Permutation tests are applied to the out-of-bag sample to generate $p$-values assessing statistical significance for predictors used within the random forest algorithm while controlling for the remaining predictors. Each variable, $W$, is selected and the values of $W$ in the out-of-bag sample are permuted $d$ times, to obtain $\mathscr{W}_{s}$, $s=1,2,...d$. This breaks their association with $Y(t)$. The out-of-bag samples corresponding to both $\mathscr{W}_{s}$ and $\mathit{W}$ are then passed through the forest to obtain new predictions, $\mathcal{P}_{\mathscr{W},s}$ and $\mathcal{P}_{\mathit{W}}$ respectively. Exploiting the fact that ${p}_{w,\text{test}}=\dfrac{\frac{1}{d}\sum_{s=1}^{d}(\mathcal{P}_{W}-\mathcal{P}_{\mathscr{W},s})}{\sigma_{\mathscr{W},\mathit{W}}}$ has an asymptotically standard normal distribution, one can then obtain a Wald-type test of variable significance ($\sigma_{\mathscr{W},\mathit{W}}$ is estimated from the permutation distribution of prediction errors).

### v3

@@ 2 Notation and Censored Longitudinal Data Structure
Figure 1 uses three hypothetical subjects to illustrate the above two steps of converting traditional recurrent event data $\{X^{*}_{i,j},\delta_{i,j}\}$ into the censored longitudinal data format $\{X_{i}(t),\delta_{i}(t):t\in\mathcal{T}_{i}\}$. Our prediction target can now be formally expressed as $P(T_{i}(t)\geq\tau|\bm{Z}_{i}(t))$, i.e., at a pre-specified time point $t$, the probability of remaining recurrent-event-free over a subsequent follow-up window of length $\tau$ given $\bm{Z}_{i}(t)$, a set of time-invariant or time-varying predictors observed up to time $t$. In the following, we will occasionally suppress the subscript $i$ and write $P(T(t)\geq\tau|\bm{Z}(t))$ for ease of presentation.
2.1 Constructing Historical Recurrent Event Information as Predictors
An important time-dependent covariate available in longitudinal studies is past recurrent event history, particularly when recurrent events within a subject are highly correlated. However, it is not always the case that history covariates are fully known at the beginning of study follow-up. Thus we consider two different types of history covariates (1) full history covariates, where $\bm{Z}_{i}(t)$ is known for each follow-up window starting at $t_{0}\in\mathcal{T}_{i}$ or (2) partial history covariates where $\bm{Z}_{i}(t)$ is unavailable for analysis at $t_{0}$, but recorded for all $t>t_{0}$. In setting (2), we perform a multiple imputation analysis where, for each imputed dataset, $m=1,\dots,M$, random variables of interest prior to $t_{0}$ for individual $i$ are independently sampled from an appropriate distribution, controlled by parameters that are themselves sampled from uniform distributions consistent with clinician knowledge of recurrent event behavior in the population.
For each follow-up window starting at $t\in\mathcal{T}_{i}$, we define predictors based on the history recurrent events which often can be predictive of future events. Three definitions are presented below based on the motivating application in this paper, though others are possible.
First, consider a dynamic predictor that averages the observed $\tau^{*}$-restricted event times at the check-in time points prior to time $t$. Specifically, in the history before $t$, identify those check-in time points: $\mathcal{T}^{*}_{i}(t)=\{t^{*}\in\mathcal{T}_{i}\cup\mathcal{T}^{0}_{i}:t^{*}+\tau^{*}\leq t\}$, where $\tau^{*}$ may differ from the original $\tau$. Here $\mathcal{T}^{0}_{i}$ is a set of check-in time points before time $t_{0}$ that are evenly spaced with $a^{*}$ time units away; it is introduced here solely for constructing predictors based on history information prior to time $t_{0}$. Here $a^{*}$ may differ from $a$ used in the construction of the censored longitudinal data structure at the beginning of Section 2. The $\tau^{*}$-restricted event time is then $\min\{T_{i}(t^{*}),\tau^{*}\}$. We average them over all the $t^{*}$ in $\mathcal{T}_{i}^{*}(t)$ to construct:
$H_{f,i}^{(1)}(t)=\left.{\sum_{t^{*}\in\mathcal{T}^{*}_{i}(t)}\min\{T_{i}(t^{*}),\tau^{*}\}}\right/{|\mathcal{T}^{*}_{i}(t)|}.$ (1)
Note that in Equation (1), at $t=t_{0}$, the definition assumes there exists historical recurrent event information for analysis. Otherwise, patient history prior to $t_{0}$ can be supplemented via (multiple) imputations of recurrent event times, $\tilde{T}_{i,j}$, which are then converted to imputes of the censored longitudinal random variable, denoted as $\widetilde{T}_{i}(t^{*})=\tilde{T}_{i,j}-t^{*}$ for $t^{*}\in\mathcal{T}^{*}_{i}(t_{0})$. We now can define a partially observed version of Equation (1):
$H_{p,i}^{(1)}(t)=\left.{\sum_{t^{*}\in\mathcal{T}^{*}_{i}(t)}I(t^{*}\leq t_{0})\times\min\{\tilde{T}_{i}(t^{*}),\tau^{*}\}+I(t^{*}>t_{0})\times\min\{T_{i}(t^{*}),\tau^{*}\}}\right/{|\mathcal{T}^{*}_{i}(t)|}.$ (2)
Second, at check-in time $t$, we may construct a predictor based on how long in the past $\tau^{*}$ period did the most recent event occur, if it exists. With the full history recurrent event information available, the time since the most recent exacerbation at $t$ is
$H^{\dagger}_{f,i}(t)=\begin{cases}\min\limits_{j:T_{i,j}<t}\{|t-T_{i,j}|\}&\text{if }\{j:T_{i,j}<t\}\neq\emptyset;\\ \tau^{*}&\text{if }\{j:T_{i,j}<t\}=\emptyset.\end{cases}$
This information can be coded as a continuous history variable, $\min\{\tau^{*},H^{\dagger}_{f,i}(t)\}$, or be coded as a categorical variable, for instance:
$H_{f,i}^{(2)}(t)=$ $\ 4\times I\left\{0\mbox{ days}\leq H^{\dagger}_{f,i}(t)<31\mbox{days}\right\}+3\times I\left\{31\mbox{ days}\leq H^{\dagger}_{f,i}(t)<93\mbox{ days}\right\}+$
$\ 2\times I\left\{93\mbox{ days}^{*}\leq H^{\dagger}_{f,i}(t)<182\mbox{ days}\right\}+I\left\{182\mbox{ days}\leq H^{\dagger}_{f,i}(t)<\tau^{*}\right\},$ (3)
where increasing levels of $H_{f,i}^{(2)}(t)$ indicate shorter periods without recurrent events in the past, and $H_{f,i}^{(2)}(t)=0$ when $H^{\dagger}_{f,i}(t)\geq\tau^{*}$. When the full history of recurrent information is not available for analysis at $t_{0}$, one may impute $\tilde{T}_{i,j}$ and identify the most recent imputed event time before $t_{0}$ as $\max\tilde{T}_{i}$. We then define a partial history version of Equation (3) as:
$H^{\dagger}_{p,i}(t)=\begin{cases}|t-\max\tilde{T}_{i}|&\text{if }t\leq t_{0};\\ |t-\max\tilde{T}_{i}|&\text{if }t>t_{0}\text{ and }\{j:T_{i,j}<t\}=\emptyset;~{}\textrm{and}{}\\ \min\limits_{j:T_{i,j}<t}\{|t-T_{i,j}|\}&\text{if }t>t_{0}\text{ and, }\{j:T_{i,j}<t\}\neq\emptyset.\\ \end{cases}$
Again, similar to the $H^{(2)}_{f,i}(t)$ case, we may code this variable as the continuous version $H_{p,i}^{(2)}(t)=\min\{\tau^{*},H^{\dagger}_{p,i}(t)\}$ or use a categorical form, as in Equation (3), replacing $H^{\dagger}_{f,i}(t)$ with $H^{\dagger}_{p,i}(t)$.
Third, the history recurrent event information predictor can be based on the frequency with which a recurrent event occurred within $\tau^{*}$ time units at each of the past check-in time points $t^{*}\in\mathcal{T}^{*}_{i}(t)$. If all the historical recurrent event information is available for analysis, we define
$H^{(3)}_{f,i}(t)=\left.{\sum_{t^{*}\in\mathcal{T}^{*}_{i}(t)}I\{{T}_{i}(t^{*})\leq\tau^{*}\}}\right/{|\mathcal{T}^{*}_{i}(t)|}.$ (4)
In the case where events prior to time $t_{0}$ are imputed as described above, we define
$H^{(3)}_{p,i}(t)=\left.{\sum_{t^{*}\in\mathcal{T}^{*}_{i}(t)}[I(t^{*}\leq t_{0})\times I\{\widetilde{T}_{i}(t^{*})<\tau^{*}\}+I(t^{*}>t_{0})\times I\{T_{i}(t^{*})<\tau^{*}\}]}\right/{|\mathcal{T}^{*}_{i}(t)|}.$ (5)
Relatedly, analysis using multiply-imputed partial history variables included in the set of dynamic predictors $\bm{Z}_{i}(t)$ requires combination via Rubin’s rule, where appropriate. Briefly, Rubin’s rule allows an analyst to combine estimates of a parameter of interest, say $\nu$, from separate analyses of imputed datasets, summarizing results with an overall point estimate and corresponding variance estimate (Rubin, 1987). Let $m=1,2,\ldots,M$ denote the number of complete (potentially imputed) datasets with point estimates $\widehat{\nu}_{m}$, and estimated variance $\widehat{\sigma}_{m}^{2}$. The combined point estimate is $\bar{\nu}=\frac{1}{M}\sum_{m=1}^{M}\widehat{\nu}_{m}$. The pooled variance, $V_{t}$, associated with $\bar{\nu}$ has two components: the within- and the between-imputation variance. We define $V_{t}=\frac{1}{M}\sum_{m=1}^{M}\widehat{\sigma}_{m}^{2}+\frac{(M+1)}{M(M-1)}\sum_{m=1}^{M}(\widehat{\nu}_{m}-\bar{\nu})^{2}$. Relying on theory of asymptotic tests for large $n$, $\frac{\bar{\nu}^{2}}{V_{t}}\sim\chi^{2}_{(1)}$. For more details, see Rubin (1996).
3 Pseudo-Observations for Dynamic Prediction
In this section, we (1) briefly review the heuristics behind the use of pseudo-observations in censored data settings and (2) develop pseudo-observations appropriate for dynamic prediction based on the censored longitudinal data structure described in Section 2.
Pseudo-observations are a variation of jack-knife methodology applicable to censored time-to-event outcomes (Andersen, Klein and Rosthøj, 2003). This technical device has been used successfully for a single time-to-event and recurrent events in manuscripts such as: Andrei and Murray (2007); Hengelbrock and others (2016); Xia, Murray and Tayob (2020); Zhao and others (2020). Essentially a pseudo-observation of a censored outcome, once created, can be used in regression contexts as if it was an uncensored outcome.
The theoretical justification behind the use of pseudo-observations typically centers on an estimand of interest for each individual $i$, which in our dynamic prediction setting is $P\{T_{i}(t)\geq\tau|\bm{Z}_{i}(t)\},$ i.e., the conditional probability of being event free in the follow-up window of length $\tau$ starting at $t$ given an individual’s covariates, $\bm{Z}_{i}(t),$ known at the beginning of this follow-up window for $i=1,\ldots,n(t)$, where $n(t)$ is the number of individuals at risk at time $t$. This conditional probability can be equivalently expressed as a conditional expectation, $E[I\{T_{i}(t)\geq\tau\}|\bm{Z}_{i}(t)]$, where $I\{T_{i}(t)\geq\tau\}$ given $\bm{Z}_{i}(t)$ is a Bernoulli$[P\{T_{i}(t)\geq\tau|\bm{Z}_{i}(t)\}]$ random variable that is not completely observed if $T_{i}(t)$ is censored prior to $\tau$, and $E[I\{T_{i}(t)\geq\tau\}|\bm{Z}_{i}(t)]$ is a random variable that depends on $\bm{Z}_{i}(t)$ and its corresponding probability distribution. The law of total expectation states that the mean of $E[I\{T_{i}(t)\geq\tau\}|\bm{Z}_{i}(t)]$ taken across the distribution of $\bm{Z}(t)$ is algebraically equivalent to $E[I\{T(t)\geq\tau\}]=P\{T(t)\geq\tau\}$, which is the (unconditional, marginal) expected value of $I\{T(t)\geq\tau\}$ in the absence of knowledge of $\bm{Z}(t)$. That is,
$P\{T(t)\geq\tau\}=E[P\{T(t)\geq\tau|\bm{Z}(t)\}],$ (6)
where the outermost expectation on the right hand side of this expression is taken with respect to the population distribution of $\bm{Z}(t)$.
The intuition behind creating a pseudo-observation for individual $i$ comes from applying Equation (6) to two different empirically observed distributions of $\bm{Z}(t)$ based on the cohort of $n(t)$ individuals contributing data during the follow-up window starting at $t$: (1) the empirically observed distribution of $\bm{Z}(t)$ based on the entire cohort, $j=1,\ldots,n(t),$ and (2) the empirically observed distribution of $\bm{Z}(t)$ based on this cohort, excluding individual $i$. For cohort (1), Equation(6) becomes
$P\{T(t)\geq\tau\}=\frac{1}{n(t)}\sum_{j=1}^{n(t)}P\{T_{j}(t)\geq\tau|\bm{Z}_{j}(t)\}.$ (7)
For cohort (2), Equation(6) becomes
$P^{(-i)}\{T(t)\geq\tau\}=\frac{1}{n(t)-1}\sum_{j\neq i}^{n(t)}P\{T_{j}(t)\geq\tau|\bm{Z}_{j}(t)\},$ (8)
where we use the notation $P^{(-i)}\{T(t)\geq\tau\}$ to indicate that individual $i$ being excluded from the empirical distribution of $\bm{Z}(t)$ affects quantities on both sides of this expression. Taking Equations (7) and (8), together and solving for $P\{T_{i}(t)\geq\tau|\bm{Z}_{i}(t)\}$, we get
$P\{T_{i}(t)\geq\tau|\bm{Z}_{i}(t)\}=n(t)P\{T(t)\geq\tau\}-\{n(t)-1\}P^{(-i)}\{T(t)\geq\tau\},$ (9)
where the left hand side of this expression is the dynamic prediction probability for patient $i$ that we would like to obtain and the right-hand side involves marginal population quantities. The right-hand side of this Equation can be estimated completely nonparametrically using Kaplan-Meier estimates $\widehat{P}(T(t)\geq\tau)$ for $P(T(t)\geq\tau)$, and $\widehat{P}^{(-i)}(T(t)\geq\tau)$ for $P^{(-i)}(T(t)\geq\tau)$; the formulae are provided in Supplemental Materials Section A1. An appropriate pseudo-observation for dynamic prediction in the recurrent event setting is one that has approximately the same mean as $I[T_{i}(t)\geq\tau]$ given $\bm{Z}_{i}(t)$, the uncensored random variable of interest. The pseudo-observation,
$\widehat{S}^{\tau}_{i}(t)=n(t)\widehat{P}(T(t)\geq\tau)-[n(t)-1]\widehat{P}^{(-i)}(T(t)\geq\tau),$ (10)
has asymptotic mean, $P[T_{i}(t)\geq\tau|\bm{Z}_{i}(t)],$ which meets this criterion. In fact, in the special case with no censoring, all pseudo-observations, $\widehat{S}^{\tau}_{i}(t),$ reduce algebraically to $I(T_{i}(t)\geq\tau),i=1,\ldots,n(t).$
4 Semiparametric Model for Dynamic Prediction (Modified XMT Model)
Pseudo-observations, $\widehat{S}^{\tau}_{i}(t),t\in\mathcal{T}_{i},i=1,\ldots,n(t)$, may be used as observed outcomes in a variety of analyses typically reserved for uncensored data, including Generalized Estimating Equation (GEE) models and longitudinal random forest algorithms. Consider the marginal model (estimated by GEE),
$\text{logit}[\widehat{S}_{i}^{\tau}(t)]=\beta^{\top}\bm{Z}_{i}(t),$ (11)
where outcomes for individual $i$, $\{\widehat{S}_{i}^{\tau}(t),\mbox{ for }t\in\mathcal{T}_{i}\}$ are assumed to have an unstructured correlation matrix. Because the originally published XMT model used pseudo-observations for $\tau$-restricted means and not our custom made pseudo-observations for dynamically predicting recurrent event probabilities, $\widehat{S}^{\tau}_{i}(t),i=1,\ldots,n(t)$, we call model (11) “the modified XMT (m-XMT) model” in this manuscript. Existing software can be used to fit this model. Estimated dynamic prediction probabilities from this model are $\widehat{P}[T_{i}(t)\geq\tau|\bm{Z}_{i}(t)]=\textrm{logit}^{-1}\{\widehat{\beta}^{\top}\bm{Z}_{i}(t)\}$. This proposed regression model is a semiparametric model for dynamic prediction of recurrent event probabilities, subject to the same limitations described for parametric and semiparametric models in the introduction; that is, difficulty in the regression specification when there are many possibly highly correlated predictors.
5 Random Forest for Dynamic Prediction (RFRE.PO algorithm)
@@ 4 Semiparametric Model for Dynamic Prediction (Modified XMT Model)
The number of possible trees is combinatorially large, challenging efficient global optimization. Greedy algorithms have been developed to approximate the optimal global tree by myopically optimizing prediction error at the start of each branch. We will focus on binary trees in this paper for their popularity and effectiveness. The loss associated with the prediction error for a branch is often termed “impurity” which measures how similarly the observations behave on either side of the binary split. The branching procedure halts, for example, when the number of observations in a terminal node or the number of terminal nodes reaches respective thresholds. Advantages of a regression tree include invariance to monotonic transformation of predictors, flexible approximation to potentially severe nonlinearities, and the capacity to approximate $L-1$ way interactions for a tree with depth $L$.
@@
Let $[\widehat{S}_{i}^{\tau}(t)$, $\bm{Z}_{i}(t)]$ be the data for individual, $i$, pertaining to the $\tau$-duration follow-up window starting at $t$, where $\widehat{S}_{i}^{\tau}(t)$ is a pseudo-observation as defined in Section 3 and $\bm{Z}_{i}(t)$ is a set of $p$-dimensional covariates available at time $t\in\mathcal{T}_{i}$ for $i=1,2,...,n$; here we use ${\mathcal{T}}_{i}$’s introduced in Section 2 with shared measurement timings to align more closely with our construction in Section 3, although the algorithm presented herein is applicable to scenarios with subject-specific measurement timings. When pseudo-observations of probability of time-to-first-recurrent-events are applied to random forests for prediction of the structure described here, we call the algorithm “RFRE.PO” (short for “Random Forest for Recurrent Events based on Pseudo-Observations”); we depict the analysis pipeline in Figure 2. Also note that the $\bm{Z}_{i}(t)$ may include time $t$. The following algorithm assumes $\widehat{S}_{i}^{\tau}(t)=g\{\bm{Z}_{i}(t)\}+\epsilon_{i}(t)$ with $\mathbb{E}[\epsilon_{i}(t)]=0$ and is designed to estimate $g(\cdot)$ which describes a potentially complex relationship between the outcome and the predictors.
@@
      (Choose branching): Choosing a covariate among $\bm{W}$ and the split value $w_{kl}$ will result in two daughter nodes, $\{N_{w_{kl},1},N_{w_{kl},0}\}$. Criteria for selecting the best variable and split vary by outcome type and metric of prediction accuracy. For continuous outcomes, the split with minimal sum of squared errors (SSE), is selected, though for other responses types, loss functions such as the Gini Index may be selected. For any particular node split under consideration using covariate $w$, that partitions individuals into two groups ($\{N_{w,0},N_{w,1}\}$) of sizes $n_{w,0}$ and $n_{w,1}$, we minimize the criterion $SSE(w)=\sum_{c=0,1}\sum_{q=1}^{n_{w,c}}(\widehat{S}_{q}^{\tau}(t)-\overline{{S}_{c}})^{2}$, where $\overline{{S}_{c}}=\frac{1}{n_{w,c}}\sum_{q=1}^{n_{w,c}}\widehat{S}_{q,c}^{\tau}(t)$, where $c=0,1$.
@@
      (Stopping criteria and prediction from the grown tree) Repeat Step 2a and 2b until one of the following hyperparameters has been met: minimum node size ($\min\{n_{w,c},c=0,1\}$) is set to $40$ in this paper) is reached, or $SSE(w)$ does not decrease any further with more branches. This results in a single binary decision tree. For each terminal node (“leaf”) of the tree $C$, predict $P\{T_{i}(t)\geq\tau\}$ by the average of the outcome for observations in leaf $C$. Let the prediction be $\tilde{P}_{b}(\bm{Z}_{i}(t);\widehat{\bm{\theta}}_{b})$, where $\tilde{\bm{\theta}}_{b}$ is the collection of terminal node predictions in the $b$-th tree, and $\bm{Z}_{i}(t)$ is the covariate profile for subject $i$.
@@
    (Ensemble prediction): The final prediction for an individual $i^{\prime}$ with covariates $\bm{Z}_{i^{\prime}}(t)$ is given by the average of all the outputs from $B$ binary decision trees above: $\tilde{P}_{i^{\prime}}\{T_{i^{\prime}}(t)\geq\tau|Z_{i^{\prime}}(t)\}=\frac{1}{B}\sum_{b=1}^{B}\tilde{P}_{b}(\bm{Z}_{i^{\prime}}(t);\widehat{\bm{\theta}}_{b})$.
@@
Permutation tests are applied to the out-of-bag sample to generate $p$-values assessing statistical significance for predictors used within the random forest algorithm while controlling for the remaining predictors. Each variable, $W$, is selected and the values of $W$ in the out-of-bag sample are permuted $d$ times, to obtain $\mathscr{W}_{s}$, $s=1,2,...d$. This breaks their association with $\widehat{S}_{i}^{\tau}(t)$. The out-of-bag samples corresponding to both $\mathscr{W}_{s}$ and $\mathit{W}$ are then passed through the forest to obtain new predictions, $\mathcal{P}_{\mathscr{W},s}$ and $\mathcal{P}_{\mathit{W}}$ respectively. Exploiting the fact that ${p}_{w,\text{test}}=\dfrac{\frac{1}{d}\sum_{s=1}^{d}(\mathcal{P}_{W}-\mathcal{P}_{\mathscr{W},s})}{\sigma_{\mathscr{W},\mathit{W}}}$ has an asymptotically standard normal distribution, one can then obtain a Wald-type test of variable significance ($\sigma_{\mathscr{W},\mathit{W}}$ is estimated from the permutation distribution of prediction errors).
6 Evaluation Metrics